# ByDeWay: Boost Your multimodal LLM with DEpth prompting in a training-free Way

### v1

@@ Abstract
We introduce a training-free framework, ByDeWay, a training-free method to boost the performances of Multimodal Large Language Models. Specifically, ByDeWay leverages a novel prompting strategy, Layered-Depth-Based Prompting (LDP), that enhances the spatial reasoning and grounding capabilities of Multimodal Large Language Models (MLLMs). Our key insight is to inject structured spatial context derived from monocular depth estimation into the input prompts—without modifying any model parameters. By segmenting scenes into closest, mid-range, and farthest depth layers and generating region-specific captions using a grounded vision-language model, we produce explicit depth-aware textual descriptions. These descriptions are concatenated with image-question prompts to guide the model toward spatially grounded and hallucination-resistant outputs. Our method is lightweight, modular, and compatible with any black-box MLLM. Evaluations on hallucination-sensitive (POPE) and reasoning-intensive (GQA) tasks show consistent improvements across multiple MLLMs, demonstrating the effectiveness of depth-aware prompting in a zero-training setup.
@@ 1 Introduction
However, persistent challenges remain, chiefly hallucination—the tendency of MLLMs to generate content that is not grounded in the visual input—and reasoning-sensitivity, where small changes in prompts or visuals can dramatically affect output. These issues become especially problematic in domains requiring precise spatial reasoning and object-level grounding, such as robotics, healthcare, and safety-critical applications. A recent survey finds that visual hallucinations—like misidentifying colors, objects, or spatial relations—continue to impede reliable deployment of MLLMs in real-world settings [4].
@@
LDP explicitly incorporates depth-aware textual context into the prompt by describing discrete spatial layers of the scene—foreground, mid-range, and background—based on captions generated from monocular depth maps. These layer-specific descriptions act as structured, interpretable anchors, capturing pseudo-3D scene organization without any model retraining.
Concretely, we use an off-the-shelf monocular depth estimator to segment input images into depth-based regions. Each region is captioned using a grounded vision–language model, yielding region-specific textual summaries. These summaries are concatenated with the original image–question pair to form a depth-enriched prompt, providing the MLLM with a clearer, spatially organized scene representation—thereby reducing hallucinations and bolstering visual reasoning in a fully training-free manner (see Figure 1).
@@ 2.2 Depth-Aware Caption Generation and Hybrid Multimodal Systems
Traditional image captioning models such as “Show and Tell” [15] and bottom-up attention [3] primarily rely on 2D visual features, overlooking depth and spatial reasoning. Recently, several works have demonstrated that incorporating depth information can significantly enhance spatial understanding in captions. For example, DEVICE [21] enriches the TextCaps task by integrating depth maps and visual-concept tokens via a depth-enhanced feature updating module, leading to more accurate relations between scene objects. Similarly, Ahmed et al. [1] propose a Transformer-based RGB–D captioning framework that explores pixel-, feature-, and hybrid-level fusion strategies on the NYU-v2 dataset, showing improved multi-sentence descriptions by leveraging depth cues.
In parallel, dense captioning approaches operating on 3D point clouds go further by explicitly segmenting objects and reasoning about their inter-object relations. Scan2Cap introduced context-aware scene-object graphs and relational attention for dense descriptions of RGB–D scans [5], while SpaCap3D [16] improved on this by introducing a token-to-token spatiality-guided Transformer encoder that learns fine-grained 3D spatial relations among objects. X-Trans2Cap [24] further extends these ideas via cross-modal knowledge transfer, distilling RGB-aware features into a point-cloud-only model for efficient inference.
@@
Our work draws on these strands by explicitly segmenting scenes according to depth and seamlessly integrating a depth-informed segmentation module into a hybrid captioning pipeline. In contrast to prior 2D-only or loosely coupled RGB–D models, we introduce a unified framework that segments the scene along depth discontinuities and delivers captions grounded in 3D spatial organization. This contributes a novel fusion mechanism that bridges the gap between depth-aware segmentation and captioning."
@@ 3 Methodology
    Region-Aware Captioning via KOSMOS-2: Each segmented region is passed individually to KOSMOS-2 [10], a grounded multimodal large language model. The model generates a localized caption describing only the visible content within the masked area. This produces three spatially grounded textual descriptions—one per depth layer.
@@ 4.2 Datasets and Task Overview
  • POPE: This benchmark comprises image-question pairs specifically curated to evaluate hallucination tendencies in MLLMs. Each sample poses a binary ("yes"/"no") question where hallucinations often arise from references to nonexistent objects. For our study, we use a subset of 150 examples and augment the prompts with structured, depth-layered captions generated through the LDP pipeline to inject explicit spatial grounding.
@@ 4.4 Metrics
$Method$ Accuracy Precision Recall F1 Score
@@ 5 Conclusions and Discussions
We introduced Layered-Depth-Based Prompting (LDP), a training-free technique for enhancing the spatial grounding capabilities of Multimodal Large Language Models (MLLMs). By injecting structured depth-aware textual descriptions—derived from monocular depth estimation and region-specific captioning—into the input prompt, LDP enables existing MLLMs to better interpret scene layout and reduce hallucination without modifying any model parameters. Experimental results show that LDP boosts performances of several multimodal large language models on both hallucination-sensitive and reasoning-sensititive tasks in a training-free way. Importantly, LDP is model-agnostic and modular—it can be used with any black-box MLLM that accepts image-instruction inputs. This makes it a practical and scalable tool for real-world deployment, especially in scenarios where retraining large models is infeasible.
@@
In summary, our work, termed as ByDeWay, highlights that prompt-level spatial grounding—powered by depth and layered captioning—is a powerful, training-free direction for improving the fidelity and robustness of modern multimodal systems.

### v2

@@ Abstract
We introduce a training-free framework, ByDeWay, a training-free method to boost the performances of Multimodal Large Language Models. Specifically, ByDeWay leverages a novel prompting strategy, Layered-Depth-Based Prompting (LDP), that enhances the spatial reasoning and grounding capabilities of Multimodal Large Language Models (MLLMs). Our key insight is to inject structured spatial context derived from monocular depth estimation into the input prompts—without modifying any model parameters. By segmenting scenes into closest, mid-range, and farthest depth layers and generating region-specific captions using a grounded vision-language model, we produce explicit depth-aware textual descriptions. These descriptions are concatenated with image-question prompts to guide the model toward spatially grounded and hallucination-resistant outputs. Our method is lightweight, modular, and compatible with any black-box MLLM. Evaluations on hallucination-sensitive (POPE) and reasoning-intensive (GQA) tasks show consistent improvements across multiple MLLMs, demonstrating the effectiveness of depth-aware prompting in a zero-training setup. The source code is available at https://github.com/Rajarshi12321/ByDeWay .
@@ 1 Introduction
However, persistent challenges remain, chiefly hallucination—the tendency of MLLMs to generate content that is not grounded in the visual input—and reasoning‑sensitivity, where small changes in prompts or visuals can dramatically affect output. These issues become especially problematic in domains requiring precise spatial reasoning and object‑level grounding, such as robotics, healthcare, and safety‑critical applications. A recent survey finds that visual hallucinations—like misidentifying colors, objects, or spatial relations—continue to impede reliable deployment of MLLMs in real‑world settings [4].
@@
LDP explicitly incorporates depth‑aware textual context into the prompt by describing discrete spatial layers of the scene—foreground, mid‑range, and background—based on captions generated from monocular depth maps. These layer‑specific descriptions act as structured, interpretable anchors, capturing pseudo‑3D scene organization without any model retraining.
Concretely, we use an off‑the‑shelf monocular depth estimator to segment input images into depth‑based regions. Each region is captioned using a grounded vision–language model, yielding region‑specific textual summaries. These summaries are concatenated with the original image–question pair to form a depth‑enriched prompt, providing the MLLM with a clearer, spatially organized scene representation—thereby reducing hallucinations and bolstering visual reasoning in a fully training‑free manner (see Figure 1).
@@ 2.2 Depth-Aware Caption Generation and Hybrid Multimodal Systems
Traditional image captioning models such as “Show and Tell” [15] and bottom-up attention [3] primarily rely on 2D visual features, overlooking depth and spatial reasoning. Recently, several works have demonstrated that incorporating depth information can significantly enhance spatial understanding in captions. For example, DEVICE [21] enriches the TextCaps task by integrating depth maps and visual-concept tokens via a depth-enhanced feature updating module, leading to more accurate relations between scene objects. Similarly, Ahmed et al. [1] propose a Transformer-based RGB–D captioning framework that explores pixel-, feature-, and hybrid-level fusion strategies on the NYU‑v2 dataset, showing improved multi-sentence descriptions by leveraging depth cues.
In parallel, dense captioning approaches operating on 3D point clouds go further by explicitly segmenting objects and reasoning about their inter-object relations. Scan2Cap introduced context-aware scene-object graphs and relational attention for dense descriptions of RGB–D scans [5], while SpaCap3D [16] improved on this by introducing a token-to-token spatiality-guided Transformer encoder that learns fine-grained 3D spatial relations among objects. X‑Trans2Cap [24] further extends these ideas via cross-modal knowledge transfer, distilling RGB-aware features into a point-cloud-only model for efficient inference.
@@
Our work draws on these strands by explicitly segmenting scenes according to depth and seamlessly integrating a depth-informed segmentation module into a hybrid captioning pipeline. In contrast to prior 2D-only or loosely coupled RGB–D models, we introduce a unified framework that segments the scene along depth discontinuities and delivers captions grounded in 3D spatial organization. This contributes a novel fusion mechanism that bridges the gap between depth-aware segmentation and captioning.”
@@ 3 Methodology
    Region-Aware Captioning via KOSMOS-2: Each segmented region is passed individually to KOSMOS-2 [10], a grounded multimodal large language model. The model generates a localized caption describing only the visible content within the masked area. This produces three spatially grounded textual descriptions, one per depth layer.
@@ 4.2 Datasets and Task Overview
  • POPE: This benchmark comprises image-question pairs specifically curated to evaluate hallucination tendencies in MLLMs. Each sample poses a binary (”yes”/”no”) question where hallucinations often arise from references to nonexistent objects. For our study, we use a subset of 150 examples and augment the prompts with structured, depth-layered captions generated through the LDP pipeline to inject explicit spatial grounding.
@@ 4.4 Metrics
MethodMethod Accuracy Precision Recall F1 Score
@@ 5 Conclusions and Discussions
We introduced Layered-Depth-Based Prompting (LDP), a training-free technique for enhancing the spatial grounding capabilities of Multimodal Large Language Models (MLLMs). By injecting structured depth-aware textual descriptions—derived from monocular depth estimation and region-specific captioning—into the input prompt, LDP enables existing MLLMs to better interpret scene layout and reduce hallucination without modifying any model parameters. Experimental results show that LDP boosts performances of several multimodal large language models on both hallucination-sensitive and reasoning-sensitive tasks in a training-free way, termed as ByDeWay. Importantly, LDP is model-agnostic and modular—it can be used with any black-box MLLM that accepts image-instruction inputs. This makes it a practical and scalable tool for real-world deployment, especially in scenarios where retraining large models is infeasible.
@@
To summarize, ByDeWay, our proposed framework, highlights that prompt-level spatial grounding—powered by depth and layered captioning—is a powerful, training-free direction for improving the fidelity and robustness of modern multimodal systems.